OpenStudy (anonymous):

The sum of two consecutive even integers is 234. What is the larger integer?

OpenStudy (anonymous):

n+ n +2 = 234 2n + 2 = 234 2n = 232 n = 116 n + 2 = 118 Thus, the larger integer is 118

OpenStudy (anonymous):

118